17. Dickinson Company has $11,820,000 million in assets. Currently half of these assets are financed with long-term debt at 9.1 percent and half with common stock having a par value of $8. Ms. Smith, Vice President of Finance, wishes to analyze two refinancing plans, one with more debt (D) and one with more equity (E). The company earns a return on assets before interest and taxes of 9.1 percent. The tax rate is 40 percent. Tax loss carryover provisions apply, so negative tax amounts are permissable.

Under Plan D, a $2,955,000 million long-term bond would be sold at an interest rate of 11.1 percent and 369,375 shares of stock would be purchased in the market at $8 per share and retired.

Under Plan E, 369,375 shares of stock would be sold at $8 per share and the $2,955,000 in proceeds would be used to reduce long-term debt.


a. How would each of these plans affect earnings per share? Consider the current plan and the two new plans. (Round your answers to 2 decimal places.)

Current Plan Plan D Plan E

Earnings per share ___________ _________ __________

b-1. Compute the earnings per share if return on assets fell to 4.55 percent. (Negative amounts should be indicated by a minus sign. Round your answers to 2 decimal places.)

Current Plan Plan D Plan E

Earnings per share ______ _______ _________


b-2. Which plan would be most favorable if return on assets fell to 4.55 percent? Consider the current plan and the two new plans.
  

  • Plan D

  • Current Plan

  • Plan E

b-3. Compute the earnings per share if return on assets increased to 14.1 percent. (Round your answers to 2 decimal place)

Current Plan Plan D Plan E

Earnings per share ______ _______ _______


b-4. Which plan would be most favorable if return on assets increased to 14.1 percent? Consider the current plan and the two new plans.
  

  • Plan D

  • Current Plan

  • Plan E

c-1. If the market price for common stock rose to $12 before the restructuring, compute the earnings per share. Continue to assume that $2,955,000 million in debt will be used to retire stock in Plan D and $2,955,000 million of new equity will be sold to retire debt in Plan E. Also assume that return on assets is 9.1 percent. (Round your answers to 2 decimal places.)
  

Current Plan Plan D Plan E

Earnings per share ______ _______ _______

c-2. If the market price for common stock rose to $12 before the restructuring, which plan would then be most attractive?

  • Plan E

  • Current Plan

  • Plan D
Answer #1

a)

I-Current Plan :First we find amount of debt & amount+number of equity

Total assets=  $11,820,000 million

Total debt= 50% of total debt = $5910000 million

Total equity = 50% of total debt= $5910000 million

Number of shares = Equity capital/par value of share = 5910000/8= 738750 million

Current Plan: We need Net Income available to equity shareholder

Return on asset before interest & tax is given as 9.1% = 11820000*9.1%=$1075620 million

Less Interest paid to debt holders @ 9.1% =5910000*9.1%= ($537810) million

Less Tax@ 40%=(1075620-537810)*40% ($215124)million

Net Income available to equity shareholder $322686 million

EPS= Net Income available to equity shareholder / Number of shares =322686/738750= 0.44

II-Plan D :First we find amount of debt & amount+number of equity

Total assets=  $11,820,000 million

Total debt= (50% of total debt)+new issue = 5910000+2955000= $8865000 million

Total equity =(50% of total debt)-buyback = 5910000 - (369375*8)=5910000-2955000= $2955000 million

Number of shares = Equity capital/par value of share =2955000/8=369375 million

Plan D :We need Net Income available to equity shareholder

Return on asset before interest & tax is given as 9.1% = 11820000*9.1%=$1075620 million

Less Interest paid to debt holders @ 9.1% =5910000*9.1%= ($537810) million

Less Interest paid to debt holders @ 11.1%=295500*11.1%= ($328005) million

Less Tax@ 40%=(1075620-537810-328005)*40% ($83922)million

Net Income available to equity shareholder $125883 million

EPS= Net Income available to equity shareholder / Number of shares =125883/369375= 0.34

III Plan E :First we find amount of debt & amount+number of equity

Total assets=  $11,820,000 million

Total debt= (50% of total debt)-repayment = 5910000-2955000=$2955000 million

Total equity =(50% of total debt)+fresh issue = 5910000 + (369375*8)=5910000+2955000= $8865000 million

Number of shares = Equity capital/par value of share =8865000/8=1108125 million

Plan E :We need Net Income available to equity shareholder

Return on asset before interest & tax is given as 9.1% = 11820000*9.1%=$1075620 million

Less Interest paid to debt holders @ 9.1% =2955000*9.1%= ($268905) million

Less Tax@ 40%=(1075620-268905)*40% ($322686)million

Net Income available to equity shareholder $484029 million

EPS= Net Income available to equity shareholder / Number of shares =484029/1108125= 0.04

Current Plan Plan D Plan e
Earnings per share (EPS) 0.44 0.34 0.04

b-1) If earning fell by 4.4% then Net Income available to equity shareholder would be as below:

Current Plan Plan D Plan e
Net Income available to equity shareholde 294289.60 97486.63 262021
Current Plan Plan D Plan e
Earnings per share (EPS) =Net Income available to equity shareholder / Number of shares 0.40 0.26 0.24

b-2) Current Plan - is better as the EPS is highest if compared with other 2 plans